solve if possible 3x+y+z=0
2x-y+z=0
2x+y+z=0

Found 2 solutions by Alan3354, edjones:
Answer by Alan3354(69443)
Answer by edjones(8007)

3x+y+z=0
2x-y+z=0
2x+y+z=0
.
2x-y+z=0
2x+y+z=0
---------add
4x+2z=0
.
2z=-4x
z=-2x
.
2x-y+z=0
2x-y-2x=0
-y=0
y=0
.
3x+y+z=0
3x+0-2x=0
x=0
.
z=0
.
Ed